Directions: Show all work. Underline final answer. To receive full credit, you must draw swap diagrams. Assume annual payment on bonds and swaps.

Company A can borrow yen at 9.6 percent and dollars at 8.1 percent. Company B can borrow

yen at 8.1 percent and dollars at 7.6 percent. If A would like to borrow yen and B would like to

borrow dollars. The financial intermediary charges a fee of 0.15. The gain is evenly split

between the two parties and exchange rate risk assumed by the intermediary. Design a swap.

What is company A’s yen rate leg and B’s dollar rate leg in the swap?
Answer #1
Company Dollar (%) Yen (%)
A 8.1 9.6
B 7.6 8.1

Company A has an advantage in Dollar and Company B has an advantage in Yen. If Company A borrows in Yen and Company B borrows in the dollar then through SWAP agreement they can change their liabilities in currency.

Total Gain = (9.6-8.1) -(8.1-7.6) = 1%

Charges of Financial Intermediaries = 0.15%

So net gain left = 1-0.15 = 0.85%

This gain is distribiuted equally in both company A and company B

Cash Flow

Company A

1. Pays to outside Lender = 8.1% dollar

2. Recieves form financial intermediary = 8.1% dollar

3. Pays to financial intermediary = 9.175 % Yen

Company B

1. Pays to outside Lender = 8.1% Yen

2. Receives from financial intermediary = 8.1% Yen

3. Pays to financial intermediary = 7.175 % dollar

Company A rate in Yen = 9.175%

Company B rate in dollar = 7.175%